OpenStudy (anonymous):

Show that if we replace x by [λ-(a/3)] and simplify, we end up with an equation that doesn't have a λ^2 term - that is, an equation in the form λ^3+pλ+q=0 (Given the general cubic equation with rational coefficients: x^3 + ax^2 + bx + c)

OpenStudy (anonymous):

Plug that term in for x in all those places and show what you get

OpenStudy (anonymous):

i got this when i expanded it all

OpenStudy (anonymous):

Okay now combine like terms

OpenStudy (anonymous):

but i found nothing to combine

OpenStudy (anonymous):

You should recheck your expansion then because a lot of that should clear out

OpenStudy (anonymous):

can you help me i dont know what i did wrong

OpenStudy (phi):

It looks ok, as far as you went.

OpenStudy (phi):

the 2nd and 3rd terms combine. there are others. match exponents on lambda and a

OpenStudy (anonymous):

ohhhhh wow im silly. i totally overlooked some of them. thank you!

OpenStudy (phi):

and I assume you will fix your bx (you changed a lambda into an x)

OpenStudy (anonymous):

yeah

OpenStudy (phi):

and you should write \[ -2a \lambda \frac{a}{3} \] as \[ -2 \lambda \frac{a^2}{3} \] so it's easier to find matching terms

OpenStudy (anonymous):

this is what i simplified so far... im not sure what to do about the b and the ab terms though or maybe i made a mistake before?

OpenStudy (phi):

you should group them in powers of lambda so that you match the coefficients with λ^3+pλ+q=0 in other words, p and q are going to be some ugly combinations of a and b

OpenStudy (phi):

your -a^2 λ should be divided by 3 ? also, you are missing the C term and the sign on bλ should be + ?

OpenStudy (anonymous):

thanks for those catches! i got down to this, does it seem to be correct?

OpenStudy (phi):

you want 27 C/27 and I got (3b-a^2)/3 for p

OpenStudy (phi):

though looking at it more carefully, your q is OK. (but I still would put 27C / 27 rather than have c dangling off to the side.)

OpenStudy (anonymous):

ok thank you so much for all your help!

OpenStudy (anonymous):

but how do you get (3b-a^2)/3 ... i can only get it with +

OpenStudy (anonymous):

oh never mind i got it

OpenStudy (anonymous):

Would you be able to help me with another problem that's related to this one?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

using the procedure from part 1, depress the equation x^3+6x^2+9x+4=0